# QiMeng-CRUX: Narrowing the Gap between Natural Language and Verilog via Core Refined Understanding eXpression

### v1

@@ Abstract
Figure 1: Descriptions of the Code-Completion ($C$) and Spec-to-RTL (SR) tasks in the VerilogEval-v2 benchmark convey equivalent design intent, but differ in structure. The top portion shows an example, while the bottom presents the performance of several general-purpose and domain-specific code LLMs on the whole benchmark. Results show that expression structure significantly impacts model performance, with SR-Recon consistently outperforming SR, and $C$ benefiting CodeVs under more constrained input formats.
@@ Introduction
Figure 4: We apply different processing strategies to the three categories. For RealSpec, we introduce prefixes and suffixes as well as interface degradation, which randomly drops the interface details to create informational variation; For CRUX, we mainly use LLM-based construction.
@@
$\mathit{pass}@k:=\mathbb{E}_{\text{problems}}\left[\frac{1-\binom{n-c}{k}}{\binom{n}{k}}\right$]. (2)
@@
The main results demonstrate our QiMeng-CRUX-V model achieves the best pass@1 performance across all benchmarks within both the General-Purpose Code Models and Verilog-Specific General Models, especially under challenging and realistic scenarios. It significantly outperforms GPT-4o on multiple tasks, most notably on RTLLM-v2, where it achieves a pass@1 of 63.8%, surpassing the previous SOTA by 12.8%, and outperforming Deepseek-V3, reaching performance comparable with Deepseek-R1. On VerilogEval-v2, QiMeng-CRUX-V brings substantial improvements over previous SOTA models across nearly all metrics, such as from 58.3% to 68.0% (+9.7%) in pass@1 on the $C$ task, and from 49.3% to 64.7% (+15.4%) on the SR task. These results suggest that CRUX offers significant advantages in tackling complex code generation tasks and real-world design inputs.
Plus, QiMeng-CRUX-V demonstrates robust performance across diverse task settings and specification styles. Originating from the CodeV Datasets, QiMeng-CRUX-V consistently performs better across tasks with diverse input forms and exhibits stronger robustness to specification variations than CodeV2.5. It achieves balanced performance across $C$ and SR (68.0% in $C$ and 64.7% in SR), avoiding the large gap observed in CodeV2.5 (58.3% in $C$ and 44.8% in SR). Furthermore, QiMeng-CRUX-V achieves substantial improvements across all metrics on the RTLLM benchmark, highlighting the effectiveness and robustness of the CRUX methodology.
@@ Appendix A Appendix
In addition to the results presented in the main results section, we have also conducted experiments on dataset scaling to assess how model performance improves with increasing dataset size. Specifically, we sampled datasets of different sizes from the original dataset, and the results of the tests on the benchmarks are shown in Figure 8. As shown in the plot, all three tasks, VerilogEval-v1-Human, VerilogEval-v2 Code Completion ($C$) and Spec-to-RTL (SR), exhibit a consistent upward trend as the dataset grows from 20k to 165k. Accuracy improves steadily across different difficulty levels and evaluation metrics, demonstrating that larger datasets consistently lead to better model performance.

### v2

@@ Abstract
Figure 1: Descriptions of the Code-Completion (CC) and Spec-to-RTL (SR) tasks in the VerilogEval-v2 benchmark convey equivalent design intent, but differ in structure. The top portion shows an example, while the bottom presents the performance of several general-purpose and domain-specific code LLMs on the whole benchmark. Results show that expression structure significantly impacts model performance, with SR-Recon consistently outperforming SR, and $C$ benefiting CodeVs under more constrained input formats.
@@ Introduction
Figure 4: We apply different processing pipelines to the three categories. RealSpec uses prefix/suffix augmentation and interface degradation for variation, while CRUX is constructed mainly via LLMs.
@@
$\mathit{pass}@k:=\mathbb{E}_{\text{problems}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]. (2)
@@
The main results demonstrate our QiMeng-CRUX-V model achieves the best pass@1 performance across all benchmarks within both the General-Purpose Code Models and Verilog-Specific General Models, especially under challenging and realistic scenarios. It significantly outperforms GPT-4o on multiple tasks, most notably on RTLLM-v2, where it achieves a pass@1 of 63.8%, surpassing the previous SOTA by 12.8%, and outperforming Deepseek-V3, reaching performance comparable with Deepseek-R1. On VerilogEval-v2, QiMeng-CRUX-V brings substantial improvements over previous SOTA models across nearly all metrics, such as from 58.3% to 68.0% (+9.7%) in pass@1 on the CC task, and from 49.3% to 64.7% (+15.4%) on the SR task. These results suggest that CRUX offers significant advantages in tackling complex code generation tasks and real-world design inputs.
Plus, QiMeng-CRUX-V demonstrates robust performance across diverse task settings and specification styles. Originating from the CodeV Datasets, QiMeng-CRUX-V consistently performs better across tasks with diverse input forms and exhibits stronger robustness to specification variations than CodeV2.5. It achieves balanced performance across CC and SR (68.0% in CC and 64.7% in SR), avoiding the large gap observed in CodeV2.5 (58.3% in CC and 44.8% in SR). Furthermore, QiMeng-CRUX-V achieves substantial improvements across all metrics on the RTLLM benchmark, highlighting the effectiveness and robustness of the CRUX methodology.
@@ Appendix A Appendix
In addition to the results presented in the main results section, we have also conducted experiments on dataset scaling to assess how model performance improves with increasing dataset size. Specifically, we sampled datasets of different sizes from the original dataset, and the results of the tests on the benchmarks are shown in Figure 8. As shown in the plot, all three tasks, VerilogEval-v1-Human, VerilogEval-v2 Code Completion (CC) and Spec-to-RTL (SR), exhibit a consistent upward trend as the dataset grows from 20k to 165k. Accuracy improves steadily across different difficulty levels and evaluation metrics, demonstrating that larger datasets consistently lead to better model performance.